Subject: APO 3.5-m Users Committee minutes 5/23/05

From: Bruce Gillespie

Submitted: Mon, 6 Jun 2005 10:04:01 -0600

Message number: 908 (previous: 907, next: 909 up: Index)

                          APO 3.5-m Users Committee Phonecon, 5/23/05

Attending:  Suzanne Hawley, Michael Strauss, Russet McMillan, Bruce  
Gillespie, Don York, Karl Glazebrook, Bruce Balick, Jon Holtzman

Absent: Al Harper, Fred Hearty, John Bally, Rene Walterbos

Minutes taken by Bruce Gillespie


User feedback, comments from Institutional representatives:

      o Michael Strauss (Princeton):  Michael said that there is  
nothing much to report, users are pretty happy, and he recently had a  
good echelle run.  SH and BG are scheduled to visit PU next week for  
general discussions with users.

      o Jon Holtzman (NMSU): Jon reported that NMSU had its first NIC- 
FPS run, which went pretty well.

      o Don York (UChicago):  Don thinks the main UC programs are  
going OK.  The GRB program is having some problems with the reduction  
of NIC-FPS data, something related to banding in the middle of the  

=> Suzanne asked that all committee members check with their NIC-FPS  
users to see if they are reducing and using NIC-FPS data, in general.

=> BruceG said that he recently talked to Josh Frieman about  
potential telescope time conflicts between the SDSSII SNe observing  
program and GRB alerts starting this fall.  Josh and Don Lamb need to  
discuss this.  Suzanne said that she thinks that UC is already trying  
to work this out.

      o Karl Glazebrook (JHU):  Karl said "nothing really, things are  

      o Suzanne Hawley (UWashington):  Suzanne said that the UW users  
are happy, lately.  The department is planning to hold a meeting  
shortly about the 3.5-m telescope and futures issues.

Jon Holtzman added that NMSU recently was awarded an NSF grant for  
its 1-m telescope, which involves providing "community use" of the  
facility.  This could play into our futures planning for the site.


What should we do with the old, backup, 2ndary mirror?:

Suzanne mentioned that after the new top-end is completed, we will  
have the old top-end as surplus, plus the old 2ndary mirror that was  
replaced several years ago by the one currently in the telescope.  If  
anybody has ideas on how we might use the old mirror and top-end for  
specialized science purposes, let us know.  Don York pointed out that  
this is a topic that can be also addressed by the Futures Committee.


Discussion on telescope/instruments and CIF report (see below):

BruceG led a short discussion on the subject report.  There were  
comments about the impact of having a bright-time shutdown in October  
to install the new top-end, which could impact the SNe followup  
observing and other programs.  Suzanne said that by August, when the  
fall schedule will be cast, we will reconsider whether or not it  
makes sense to schedule the shutdown in October, both from the  
perspective of impact to observing, and the state of readiness of the  
new top-end project.


Goddard Fabry-Perot instrument initiatives:

Suzanne said she is pleased to see the level of involvement and  
interest by Bruce Woodgate, the GSFC PI of the Goddard Fabry-Perot  
instrument, which has been a visiting instrument at the 3.5-m  
telescope for several years.  Don York and BruceW have recently  
proposed building filter holders for the complete set of 100-Angstrom  
pre-filters that belong to the instrument, six filters per wheel,  
which could be installed quickly in the F-P for medium-bandpass  
imaging by anybody.  Also proposed is a new lower-noise detector, and  
a tip-tilt module for the NA2 port that would be usable by all NA2  
instruments.  With these and some housekeeping upgrades and a TUI  
interface, the F-P would become a powerful and easier-to-use  
instrument for the general ARC community.  BruceW is also considering  
building a higher resolution NIR spectrograph (R>9000) that would use  
a zero-read-noise chip, sensitive in the .8 to 1.6 micron spectral  
range and may be interested in bringing it to APO.


Status of UVa discussions:

Suzanne said that UVa has returned to earnest discussions with us  
regarding ARC membership, and in providing a new NIR spectrograph  
(aka Triplespec) for the telescope.  Also, BruceG mentioned that UVa  
has asked if the CorMASS low-res NIR spectrograph could be returned  
to APO as a visiting instrument this summer, for use on the 3.5-m  
telescope in the fall and winter.  Michael asked what the limiting  
magnitude of the instrument is on our telescope, and BruceG said that  
it's about 15th magnitude in K in a half-hour exposure.  Suzanne  
asked the committee members to poll their users on their potential  
interest in using CorMASS in Q4; in the meantime we will tell UVa  
that they can make plans to have the instrument shipped to APO.


A new DIS red detector?:

Jon Holtzman described a proposal to replace the DIS red detector  
with a pin-compatible deep-depletion device, which would, in  
principle, mitigate the red fringing problem while not compromising  
noise performance or quantum efficiency.  The published specs on a  
chip from e2v (the same mfg. as for the existing chip we have) are  
very promising, and the modest funding required for purchase has been  
identified from an external source.  There was a group discussion  
about whether anybody knows if deep-depletion devices have been  
successfully used in astronomy.  Karl knows of similar devices in the  
field that do have lower fringing, decent read noise, but they seem  
to be somewhat more susceptible to cosmic ray events.  Suzanne talked  
to Chris Stubbs who thought they were in use in Megacam at the MMT;  
ask Brian McLeod at CfA.  Someone said they also may be used in  
GMOS.  Jon Holtzman agreed to ask e2v if they could tell him who has  
used their deep-depletion chips.

Karl asked if Jon Holtzman could send him the chip specifications,  
and Jon agreed to do this.  Don York said that we should ask Josh  
Frieman what chips they are using in the Dark-energy Camera.  Suzanne  
pointed out that if/when we replace the red chip in DIS, some effort  
will be made to identify and fix the intermittent"red-noise" problem  
we have been having.


Recent telescope scattered light results:

Jon Holtzman said that recent measurements of the stray light in the  
3.5-m telescope, after completing the baffles improvements, indicate  
that we can now do point-source photometry with the telescope, to a  
precision of ~1%.  This is a happy ending to a long story.


Summer shutdowns schedule:

Susanne said that the summer shutdowns are now scheduled, all in  
bright time:

      o 18 - 29 July;

      o 15 - 26 August; and

      o 27 - 28 August is NIC-FPS engineering time


Plan for phasing-out Remark:

Suzanne said that the only Remark capabilities that are not done by  
TUI are the echelle guiding and instrument control computer; these  
should be TUI-ized by the end of June.  We expect to see support for  
Remark phased out during June, and no Remark usage in Q3.  The small  
number of Remark holdouts will be coaxed to move to TUI.


GRB policies baselined, GRB documents website, general discussion:

BruceG briefly mentioned that APO is establishing a GRB website (at a  
blind URL) where various GRB documents and e-mail will be archived  
and available to the ARC users.  When completed, an announcement will  
be made to the ARC GRB team and to the Users Committee members.


Exo-planet Tracker prototyping and prospects:

Suzanne reported that the ET team from UFlorida had good "proof-of-  
concept" runs at the SDSS telescope this spring.  Decent radial  
velocities were measured, although there were problems (with known  
causes and solutions) with system throughput; it's not a real  
"system" yet, only a prototype.  They were also able to demonstrate  
the "piggy-back" mode where ET data and SDSS spectra were taken  
simultaneously.  Jian Ge will submit a proposal to SDSS-II for a  
Phase 2 demonstration next year, and he has expressed interest in  
being more involved with ARC projects in general, including the 3.5-m  
telescope.  This could include a NIR high-res spectrograph and/or an  
ET single-object instrument.  This needs to be factored into the  
thinking of the Futures Committee.


APOLLO lunar laser ranging update:

BruceG said that Tom Murphy is ramping up the APOLLO project with  
operations likely by the end of the year.  As it gets closer to being  
a reality, we will need to look carefully at site support issues,  
telescope scheduling, and other concerns.  Bruce also pointed to a  
very nice two-page article on the APOLLO project in the July issue of  
Sky and Telescope, which was just published.


Next phonecon will be on Monday, 20 June, at 8:30 AM Pacific Time.  
Agenda and other materials will be sent to the Users Committee  
members during the preceding week.


      3.5-m Telescope, Instruments, and CIF Projects Highlights, 4/23  
through 5/17/05

                                  Bruce Gillespie

0) Overview

The telescope and instruments generally performed well during the  
period, with typical spring-like  periods of good observing weather  
and some excellent seeing.  A handful of telescope and instrument  
problems were addressed, or are being worked on, as given below.  A  
small amount of usable observing time was lost to these problems,  
less than the equivalent of a night.

1) Telescope

      o cleaned optics twice (this is the "dusty" season)

      o cleaned drive surfaces (ditto)

      o enclosure drive/servo problems:

          During the past month there have been recurring problems  
with the enclosure drive systems.  At the present time, the enclosure  
is working fine but testing and review of telemetry has not indicated  
what the underlying cause of the intermittent problem is.  Some  
needed repair work was done to the enclosure controller electronics  
after the first occasion, but the problem seemed to return after more  
than a week later, then disappeared, and then happened again this  
past weekend.  Much effort is being spent trying to debug this, but  
at this time the telescope is behaving normally, which makes finding  
and fixing the problem difficult.  More testing and analysis is  

2) Instruments

      o echelle controller problems; repairs, upgrades:

          We have had infrequent, but long-standing, problems with  
data transfers between the echelle and its Instrument Control  
Computer (ICC).  We tried a fix which inadvertently resulted in the  
total dysfunction of the data acquisition system for the instrument.  
We were eventually able to fix this problem by replacing several  
aging components, but the intermittent problem is probably still with  
us.  This made us aware of the increasing vulnerability we have in  
using older computers in operationally critical places, so a decision  
was taken to upgrade the ICC.  Working with the Dale Sandford at  
Yerkes, an ICC replacement project has been started that will replace  
the ICC and s/w with a system that uses Linux, is largely the same as  
the DIS s/w, is TUI-friendly, and undoubtedly buys us increased  
reliability and maintainability.  We expect this ICC upgrade to be  
completed in a matter of a few weeks.

      o GRIM2 stored, plan for surplus:

          Al Harper expressed interest in recovering GRIM2 for  
scientific and teaching purposes, so we have tentatively agreed to  
have him take the instrument to Yerkes.  We are discussing with Al  
the possibility of having the GRIM2 grism removed for possible use in  
NIC-FPS, and returning the special filters used for planetary  
programs to their owner.

      o NIC-FPS degassed, ion gauge failure:

          NIC-FPS was degassed and vacuum-pumped, which is currently  
necessary on a ~monthly basis.  No observing time was lost, and this  
problem is expected to be fixed during the summer shutdowns.  Also,  
the ion pressure gauge failed, and a temporary repair was made.

      o GFP run:

          The Goddard Fabry-Perot was successfully used for a science  
run, and several discussions were held regarding possible upgrades to  
the instrument to make it more capable and available as a more  
general-use, facility-class instrument.

      o DIS noise:

          The pattern-noise problem in the red camera came and went  
much as it has in past months, with no real change in frequency or  
severity.  Suspicion centers on the problem being in the dewar  
electronics.  Jon Holtzman is working on a plan to attack the  
problem, as well as the problem of fringing seen in the red and near-  
IR data.

3) Engineering and CIF projects

      o Top-end CDR:

          The Critical Design Review for the new top end was held at  
APO 11-12 May, and was largely successful.  Final design work and  
fabrication will proceed towards an installation on the telescope  
during the October bright time.  There appear to be no significant  
technical nor cost risks.  An extensive set of telescope image motion  
data was presented during review, which provides a solid performance  
baseline for the current top end to compare against when the new top  
end is installed.

      o TUI:

          Work continued to enable full TUI guiding, SPIcam control,  
and DIS slit viewer control, which are the remaining parts of the TUI  
s/w that are unfinished.  Testing of TUI guiding with the echelle is  
planned in the next two weeks.

      o New drive boxes:

          Final fabrication is going fairly well, and the new boxes  
are planned for installation on the telescope during the summer  

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 908 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to